# Testing identifying assumptions in Tobit Models

### v2

@@ 6. Simulations
In this section, we provide simulation exercises for the proposed tests. The testing procedure described relies on testing sharp equalities that should hold for any arbitrary partition of the outcome, treatment and instrument support. When both the outcome and treatment are continuous, evaluating all possible equalities is technically challenging. We focus on a non-sharp set of the equalities by evaluating them at different grid partitions of their support, in a similar spirit to Honoré and Hu (2020) which is particularly well suited for continuous supports. Naturally, this choice makes the test less powerful as we don’t consider the continuum of equalities derived in Section 4, but is justified by the ease of implementation of the procedure based on intersection bounds and the performance of the test on the simulations below.
@@
When implementing the intersection bounds in STATA using the package clrbound Chernozhukov et al. (2015), the researcher must determine the range of values of the exogenous variable for which each equality will be evaluated. To guarantee the feasibility of the procedure, we adjust the evaluation points for $D$ ($Z$) to the first and $99^{th}$ percentiles of the exogenous variable in each partition of the support for the outcome $Y$ [^13] See details on implementation on the simulation replication STATA code.
@@
To consider the test’s performance under violations of the exogeneity assumption, we modify the DGP in Equation 6.8 with different values for $\rho$, reflecting various degrees of treatment endogeneity. Larger values of $\rho$ produce more acute violations of the null hypothesis. Table 2 presents the results for the Classic Tobit test: columns 3-4 report the MSE, while the remaining columns show empirical rejection rates. As expected, the power of the test increases with larger $\rho$ and bigger sample sizes. [^14] Simulations for test performance when violations of normality of the latent errors occur are presented in Appendix F.
@@
$N$ $\rho$ MSE$(\beta_{0})$ MSE$(\beta_{1})$ $\alpha=10\%$ $\alpha=5\%$ $\alpha=1\%$
5000 0.10 0.0005 0.0005 8.60% 6.00% 3.20%
0.50 0.0005 0.0104 16.03% 10.82% 5.21%
0.75 0.0005 0.0227 36.00% 27.80% 14.20%
0.80 0.0005 0.0259 42.00% 30.60% 16.80%
0.90 0.0004 0.0326 62.40% 48.40% 27.60%
0.95 0.0004 0.0360 71.34% 59.32% 35.87%
8000 0.10 0.0003 0.0005 8.40% 5.00% 2.00%
0.50 0.0003 0.0102 17.43% 12.22% 5.61%
0.75 0.0003 0.0225 45.58% 36.14% 20.08%
0.80 0.0003 0.0256 53.80% 39.80% 18.00%
0.90 0.0003 0.0323 79.00% 65.40% 36.40%
0.95 0.0002 0.0363 85.57% 75.15% 49.70%
10000 0.10 0.0003 0.0005 10.04% 7.43% 2.41%
0.50 0.0002 0.0101 15.80% 10.00% 3.80%
0.75 0.0002 0.0226 44.80% 32.80% 16.00%
0.80 0.0002 0.0256 56.60% 41.00% 21.60%
0.90 0.0002 0.0326 79.60% 66.60% 36.60%
0.95 0.0002 0.0361 90.56% 79.52% 51.61%
@@ 7.1. Alternative Approaches
There is a vast literature on alternatives to the Tobit Model that can be implemented in the presence of censored dependent variables. Most approaches consider changes or relaxations of one of the two main assumptions associated with the model. The first assumption is the parametric distribution of the error terms and latent index form connecting treatment (and covariates) to the outcome. The second assumption is the exogeneity of the treatment or potential instrument. Here, we provide a non-exhausting survey of existing work.
@@
Cragg (1971) maintains the normality of the errors, linearity of the index and exogeneity but relaxes the way censoring occurs in comparison to the latent structure of the censored outcome. Specifically, while the latent outcome is still modelled by $Y^{*}=\alpha_{0}+\alpha_{1}D+U$, they allow for the censoring to depend on a different linear index, $P(Y^{*}>0)=P(\gamma_{0}+\gamma_{1}D+e)$, increasing the model’s flexibility.
@@
Powell (1984) relaxes the parametric structure of the errors while maintaining the latent linear index and treatment exogeneity, and estimates the parameters of interest by least absolute deviations. Powell (1986) also maintains linearity of the latent index and treatment exogeneity, but relaxes normality by imposing symmetrical distributions to the latent errors, which leads to estimation by symmetrically censored least squares. Newey (1987a) relaxes exogeneity of the treatment, relying on normality and an instrumental variable to identify the model, which is estimated by generalized least squares.
@@
Honoré and Powell (1994) relax linearity and mean independence of the unobservable with respect to the treatment to exploit the idea that, although $Y^{*}_{i}-\alpha_{0}-\alpha_{1}D_{i}$ is not mean-independent of $D_{i}$, one can trim any pair of residuals $Y^{*}_{i}-\alpha_{0}-\alpha_{1}D_{i}$ and $Y^{*}_{j}-\alpha_{0}-\alpha_{1}D_{j}$, and the trimmed residuals are independent and identically distributed conditional on $D_{i},D_{j}$. They estimate the model by identically censored least absolute deviations and identically censored least squares (ICLS). Das (2002) estimates a model using symmetrically censored least squares that relaxes exogeneity of the treatment and normality of the errors. To achieve that they rely on instrumental variables, linearity of the mean of the structural error conditional on the reduced form error, and mean independence of the reduced form error.
@@
Blundell and Powell (2007) proposes a control variable approach that relaxes exogeneity and normality but maintains the latent linear structure ($\alpha_{0}+\alpha_{1}D+U$). Crucially, they impose that the distribution (or quantiles) of the latent error conditional on the treatment and instrument is only a function of the control variable $V=D-\pi(Z)$, which isolates the endogenous variation on the treatment. [^15] This assumption is weaker than independence of all errors and instruments since it does not impose $V$ independent of $Z$ but is neither stronger nor weaker than independence of $U$ and $Z$, since it permits $Z$ to affect $U$ through $V$. This allows them to estimate the effect of the treatment by censored quantile instrumental variable regression augmented by a control variable given by the quantiles of $U$ conditional on $V$ at the quantile of interest. In a similar spirit, Chernozhukov, Fernández-Val, and Kowalski (2015) focuses on conditional quantile functions and flexible approaches to estimate the control variable in the first stage.
@@
Finally, Chesher, Kim, and Rosen (2023) provides partial identification results for a general alternative by relaxing the exogeneity of the treatment and instrument, linearity of the latent index and imposing no parametric structure of the error term. They characterize the identified set for the parameters of interest following the Generalized Instrumental Variables framework (Chesher and Rosen, 2017), relying on the assumption that the relationship of $Y^{*}$ to treatment and errors is continuous and monotonic in the errors. Their approach uses the residual sets associated with the structure of the latent function and conditional probability of the error term given potential instruments.
@@ 7.2. Partial identification under monotonicity
In this subsection, we present an approach that partially identifies the effect of an endogenous treatment variable by replacing the normality and exogeneity assumptions with a monotonicity in treatment selection constraint. While less general than Chesher, Kim, and Rosen (2023), this approach is easy to implement and could be useful to empirical researchers.
@@
Assumption 6 is common in the partial identification literature (Jiang, Chiba, and VanderWeele, 2014; Manski, 1997; Manski and Pepper, 2000; Okumura and Usui, 2014). In this context, we restrict the latent selection to be monotonic with respect to the treatment. This assumption is embedded in the classic Tobit model since the inverse mills ratio, $\lambda(\cdot)$, is monotonic (and decreasing) in the treatment variable. Furthermore, independence between $D$ and $U$ restricts the sign of the coefficient of the selection term directly, as the derivative of the inverse mills-ratio is $\lambda^{\prime}(\alpha_{0}+\alpha_{1}D)\alpha_{1}$. Thus, without imposing independence or a parametric latent structure, we maintain a relevant property of the Tobit model that aids identification. Since it is not as restrictive as imposing a parametric structure and independence, we can only partially identify the parameter of interest.
Under Assumption 6 and the model described in 7.3, treatment and outcome are not independent. Note that,
@@
$E(Y|D=d,Y>0)$displaystyle-- $\alpha_{0}-\alpha_{1}d=\Gamma(d)$
@@
$\alpha_{1}$ $>$ $\frac{\partial E(Y|D=d,Y>0)}{\partial d}\text{, for all d.}$. (7.12)
@@
A bootstrap procedure could be used for a nonparametric estimator of $\frac{\partial E(Y|D=d,Y>0)}{\partial d}$ using a local polynomial regression. In this case, the estimator is asymptotically normal and converges at a nonparametric rate that depends on the bandwidth $h$. Recent developments in Calonico, Cattaneo, and Farrell (2018, 2022) for estimation and optimal coverage error bandwidth and kernel selection methods that are nonparametric robust bias-corrected (RBC) can be used through their convenient implementation using the package nprobust (Calonico, Cattaneo, and Farrell, 2019).
@@
A similar argument holds when the treatment variable is continuous. Tighter bounds are achieved by considering the maximum value of $\frac{\partial E(Y|D=d,X=x,Y>0)}{\partial d}$ across all possible points in the support for $D$ and $X$.
8. Empirical Illustration: Lee (1995)
In this section, we implement the proposed test to the data from Lee (1995). Using the 1987 cross-section of the Michigan Panel Study of Income Dynamics, the authors study the responses of married women’s labor supply ($Y$) - measured in hours per year - to hundreds of dollars in “other” household income ($D$), which is endogenous. The instrumental variables explored are dummy variables for the husband’s occupation ($Z$), which implies the restrictive identifying assumption that the wife’s labor supply is affected by the husband’s occupation only through their income. Following the original study, we add other covariates ($X$) in the linear index for both the outcome and treatment equations, controlling for factors that could impact women’s labor supply. Those include a quadratic on the person’s age, their years of completed education, the number of children coded in three categories (children up to 5 years old, ages 6 to 13, ages 14 to 17), the local unemployment rate in percentage points, and indicators for race (0 if white, 1 otherwise), homeownership (1 if owner, 0 otherwise) and if the couple has a mortgage on their home (1 if yes, 0 otherwise).
@@ 8. Empirical Illustration: Lee (1995)
Following Lee (1995), we proceed with the analysis focusing on the data for married couples with non-negative family total income or “other” household income and where the wife was of working age (18-64) and not self-employed. These selections results in 3,277 observations, for which 26 percent of wage observations are censored. Table LABEL:Table:Lee presents the estimates obtained using the IV Tobit model.
Table 4. IV Tobit specification for Lee (1995)
@@
Other household income $-0.973^{***}$
@@
The first column presents the first-stage estimates indicating the relevance of the potential $IV$. The second column presents the structural equation reflecting the effect the treatment variable has on the outcome equation. The parameter $\rho$ shows evidence of no correlation between the unobservables driving the “other” household income and hours worked after controlling for covariates. The estimated coefficient of interest indicates that other household income negatively affects the wife’s labor supply after conditional on the household characteristics. In particular, for women working positive hours, an increase of one thousand dollars in household income from other sources is estimated to reduce hours worked by 9.7 hours per year. The direction of the impacts at the intensive margin of hours worked follow intuitive patterns and are qualitative similar to those in Lee (1995). However, the model is rejected at conventional significance levels when we test for the IV-Tobit model’s validity, ($\hat{\theta}_{0.99}=0.1335>0$, $\hat{\theta}_{0.95}=0.1368>0$, and $\hat{\theta}_{0.90}=0.1385>0$). This indicates that the assumptions underlying the IV-Tobit model are not compatible with the empirical distribution of the data, and caution is needed when relying on the results.
@@
$\alpha_{1}$0.973 -0.419
conf.: confidence; LB: lower bound
The first column of Table 5 repeats the estimate for $\alpha_{1}$ from Table 4. The second column reports the estimated lower bound for $\alpha_{1}$, obtained under Assumption 6 and latent linear index only based on the nonparametric estimate of the average derivative of the conditional expectation for yearly hours worked with respect to the household income from other sources, as described in Remark 7. Even after relaxing the normality of errors and treatment exogeneity, the lower bound for the size of the effect of having higher household income from sources other than the wife’s labor on their labor supply indicates an effect larger than -4.19 hours worked per year for married women. Hence, we can rule out annual reductions of more than 4.2 hours in female labor supply for every one thousand dollars in other household income, but cannot reject that the effect is zero or positive.

### v3

@@ 6. Simulations
In this section, we provide simulation exercises for the proposed tests. The testing procedure described relies on testing sharp equalities that should hold for any arbitrary partition of the outcome, treatment and instrument support. When both the outcome and treatment are continuous, evaluating all possible equalities is technically challenging. We focus on a non-sharp set of the equalities by evaluating them at different grid partitions of their support, in a similar spirit to honore2020selection which is particularly well suited for continuous supports. Naturally, this choice makes the test less powerful as we don’t consider the continuum of equalities derived in Section 3, but is justified by the ease of implementation of the procedure based on intersection bounds and the performance of the test on the simulations below.
@@
When implementing the intersection bounds in STATA using the package clrbound chernozhukov_implementing_2015, the researcher must determine the range of values of the exogenous variable for which each equality will be evaluated. To guarantee the feasibility of the procedure, we adjust the evaluation points for $D$ ($Z$) to the first and $99^{th}$ percentiles of the exogenous variable in each partition of the support for the outcome YY. [^15] See details on implementation on the simulation replication STATA code.
@@
To consider the test’s performance under violations of the exogeneity assumption, we modify the DGP in Equation 6.8 with different values for $\rho$, reflecting various degrees of treatment endogeneity. Larger values of $\rho$ produce more acute violations of the null hypothesis. Table 2 presents the results for the Classic Tobit test. As expected, the power of the test increases with larger $\rho$ and bigger sample sizes.
@@
NN $\rho$ $\alpha=10\%$ $\alpha=5\%$ $\alpha=1\%$
5000 0.10 8.60% 6.00% 3.20%
0.50 16.03% 10.82% 5.21%
0.75 36.00% 27.80% 14.20%
0.80 42.00% 30.60% 16.80%
0.90 62.40% 48.40% 27.60%
0.95 71.34% 59.32% 35.87%
8000 0.10 8.40% 5.00% 2.00%
0.50 17.43% 12.22% 5.61%
0.75 45.58% 36.14% 20.08%
0.80 53.80% 39.80% 18.00%
0.90 79.00% 65.40% 36.40%
0.95 85.57% 75.15% 49.70%
10000 0.10 10.04% 7.43% 2.41%
0.50 15.80% 10.00% 3.80%
0.75 44.80% 32.80% 16.00%
0.80 56.60% 41.00% 21.60%
0.90 79.60% 66.60% 36.60%
0.95 90.56% 79.52% 51.61%
Based on $500$ replications.
To consider a violation with respect to the violation of the normality of errors. We consider a modified DGP from Equation 6.8, by seting $(V,Z)^{\top}\sim\mathcal{N}(\mathbf{0}_{2},\mathbf{I}_{2})$ and $U\sim F(\cdot)$. Table 3 presents the power results when $F(\cdot)$ is uniform, lognormal, and t-student distributions. As expected, the rejection rate for the t-student test with 80 degrees of freedom is close to the nominal size of the test, since it represents a very mild violation of the normality assumption. In the other three cases, which deviate significantly from the normality assumption, the test rejects the null hypothesis in most instances. This example demonstrates that the proposed test is effective in identifying violations of the distributional assumptions in the classic Tobit model.
Table 3. Classic Tobit test power for violations in the error structure
NN $U$ $\alpha=10\%$ $\alpha=5\%$ $\alpha=1\%$
5000 t-student (df=80) 7.21% 4.01% 1.40%
t-student (df=5) 92.20% 81.80% 48.80%
LogNormal 100.00% 100.00% 100.00%
Uniform 100.00% 100.00% 100.00%
8000 t-student (df=80) 7.00% 4.80% 1.20%
t-student (df=5) 99.20% 97.80% 87.58%
LogNormal 100.00% 100.00% 100.00%
Uniform 100.00% 100.00% 100.00%
10000 t-student (df=80) 5.42% 2.41% 1.00%
t-student (df=5) 99.20% 98.40% 90.58%
LogNormal 100.00% 100.00% 100.00%
Uniform 100.00% 100.00% 100.00%
@@ 7.1. Alternative Approaches
There is a vast literature on alternatives to the Tobit Model that can be implemented in the presence of censored dependent variables. Most approaches consider changes or relaxations of one of the two main assumptions associated with the model. The first assumption is the parametric distribution of the error terms and latent index form connecting treatment (and covariates) to the outcome. The second assumption is the exogeneity of the treatment or potential instrument. Here, we provide a non-exhaustive survey of existing work.
@@
cragg1971some maintains the normality of the errors, linearity of the index and exogeneity but relaxes the way censoring occurs in comparison to the latent structure of the censored outcome. Specifically, while the latent outcome is still modelled by $Y^{*}=\alpha_{0}+\alpha_{1}D+U$, they allow for the censoring to depend on a different linear index, $P(Y^{*}>0)=P(\gamma_{0}+\gamma_{1}D+e)$, increasing the model’s flexibility.
@@
POWELL1984303 relaxes the parametric structure of the errors while maintaining the latent linear index and treatment exogeneity, and estimates the parameters of interest by least absolute deviations. powell1986symmetrically also maintains linearity of the latent index and treatment exogeneity, but relaxes normality by imposing symmetrical distributions to the latent errors, which leads to estimation by symmetrically censored least squares. newey1987efficient relaxes exogeneity of the treatment, relying on normality and an instrumental variable to identify the model, which is estimated by generalized least squares.
@@
honore1994pairwise relax linearity and mean independence of the unobservable with respect to the treatment to exploit the idea that, although $Y^{*}_{i}-\alpha_{0}-\alpha_{1}D_{i}$ is not mean-independent of $D_{i}$, one can trim any pair of residuals $Y^{*}_{i}-\alpha_{0}-\alpha_{1}D_{i}$ and $Y^{*}_{j}-\alpha_{0}-\alpha_{1}D_{j}$, and the trimmed residuals are independent and identically distributed conditional on $D_{i},D_{j}$. They estimate the model by identically censored least absolute deviations and identically censored least squares (ICLS). das2002estimators estimates a model using symmetrically censored least squares that relaxes exogeneity of the treatment and normality of the errors. To achieve that they rely on instrumental variables, linearity of the mean of the structural error conditional on the reduced form error, and mean independence of the reduced form error.
@@
blundell2007censored proposes a control variable approach that relaxes exogeneity and normality but maintains the latent linear structure ($\alpha_{0}+\alpha_{1}D+U$). Crucially, they impose that the distribution (or quantiles) of the latent error conditional on the treatment and instrument is only a function of the control variable $V=D-\pi(Z)$, which isolates the endogenous variation on the treatment. [^16] This assumption is weaker than independence of all errors and instruments since it does not impose $V$ independent of $Z$ but is neither stronger nor weaker than independence of $U$ and $Z$, since it permits $Z$ to affect $U$ through $V$. This allows them to estimate the effect of the treatment by censored quantile instrumental variable regression augmented by a control variable given by the quantiles of $U$ conditional on $V$ at the quantile of interest. In a similar spirit, chernozhukov2015quantile focuses on conditional quantile functions and flexible approaches to estimate the control variable in the first stage.
@@
Finally, chesher2023iv provides partial identification results for a general alternative by relaxing the exogeneity of the treatment and instrument, linearity of the latent index and imposing no parametric structure of the error term. They characterize the identified set for the parameters of interest following the Generalized Instrumental Variables framework (chesher2017generalized), relying on the assumption that the relationship of $Y^{*}$ to treatment and errors is continuous and monotonic in the errors. Their approach uses the residual sets associated with the structure of the latent function and conditional probability of the error term given potential instruments.
@@ 7.2. Partial identification under monotonicity
In this subsection, we present an approach that partially identifies the effect of an endogenous treatment variable by replacing the normality and exogeneity assumptions with a monotonicity in treatment selection constraint. While less general than chesher2023iv, this approach is easy to implement and could be useful to empirical researchers.
@@
Assumption 6 is common in the partial identification literature (jiang2014monotone; manski1997monotone; manski2000amonotone; okumura2014concave). In this context, we restrict the latent selection to be monotonic with respect to the treatment. This assumption is embedded in the classic Tobit model since the inverse mills ratio, $\lambda(\cdot)$, is monotonic (and decreasing) in the treatment variable. Furthermore, independence between $D$ and $U$ restricts the sign of the coefficient of the selection term directly, as the derivative of the inverse mills-ratio is $\lambda^{\prime}(\alpha_{0}+\alpha_{1}D)\alpha_{1}$. Thus, without imposing independence or a parametric latent structure, we maintain a relevant property of the Tobit model that aids identification. Since it is not as restrictive as imposing a parametric structure and independence, we can only partially identify the parameter of interest.
Under Assumption 6 and the model described in equations (7.3), treatment and outcome are not independent. Note that,
@@
$E(Y|D=d,Y>0)$ - $\alpha_{0}-\alpha_{1}d=\Gamma(d)$
@@
$\alpha_{1}$ $>$ $\frac{\partial E(Y|D=d,Y>0)}{\partial d}\text{, for all d.}$ (7.12)
@@
A bootstrap procedure could be used for a nonparametric estimator of $\frac{\partial E(Y|D=d,Y>0)}{\partial d}$ using a local polynomial regression. In this case, the estimator is asymptotically normal and converges at a nonparametric rate that depends on the bandwidth $h$. Recent developments in calonico2018effect; calonico2022coverage for estimation and optimal coverage error bandwidth and kernel selection methods that are nonparametric robust bias-corrected (RBC) can be used through their convenient implementation using the package nprobust (calonico2019nprobust). Alternatively, since $\alpha_{1}>sup_{d}\frac{\partial E(Y|D=d,Y>0)}{\partial d}$, one could consider obtaining confidence regions for $\alpha_{1}$ using a CLR approach similar to that described in Section 5 by considering nonparametric estimates of $\frac{\partial E(Y|D=d,Y>0)}{\partial d}$ or its discrete counterpart. [^17] We thank an anonymous referee for this suggestion.
@@
A similar argument holds when the treatment variable is continuous. Tighter bounds are achieved by considering the maximum value of $\frac{\partial E(Y|D=d,X=x,Y>0)}{\partial d}$ across all possible points in the support for $D$ and $X$. In practice, nonparametric estimates of these derivatives can be difficult, even for moderate numbers of covariates, and particularly challenging when multiple continuous covariates are in the conditioning set. The estimated values might be unstable, in which case using the maximum estimated value can lead to unreasonable bounds for $\alpha_{1}$. An easier to implement conservative alternative is to use the average derivatives with respect to $D$, $E\left[\frac{\partial E(Y|D=d,X=x,Y>0)}{\partial d}\right]$ and rely on bootstrapped standard errors for inference.
8. Empirical Illustration: lee1995semi
In this section, we implement the proposed test to the data from lee1995semi. [^18] Data availability statement: The data that support the findings of this study are openly available in the Journal of Applied Econometrics Data Archive at http://dx.doi.org/10.15456/jae.2022313.1130270920. Using the 1987 cross-section of the Michigan Panel Study of Income Dynamics, the authors study the responses of married women’s labor supply ($Y$) - measured in hours per year - to hundreds of dollars in “other” household income ($D$), which is endogenous. The instrumental variables explored are dummy variables for the husband’s occupation ($Z$), which implies the restrictive identifying assumption that the wife’s labor supply is affected by the husband’s occupation only through their income. Following the original study, we add other covariates ($X$) in the linear index for both the outcome and treatment equations, controlling for factors that could impact women’s labor supply. Those include a quadratic on the person’s age, their years of completed education, the number of children coded in three categories (children up to 5 years old, ages 6 to 13, ages 14 to 17), the local unemployment rate in percentage points, and indicators for race (0 if white, 1 otherwise), homeownership (1 if owner, 0 otherwise) and if the couple has a mortgage on their home (1 if yes, 0 otherwise).
@@ 8. Empirical Illustration: lee1995semi
Following lee1995semi, we proceed with the analysis focusing on the data for married couples with non-negative family total income or “other” household income and where the wife was of working age (18-64) and not self-employed. These selections results in 3,277 observations, for which 26 percent of wage observations are censored. Table LABEL:Table:Lee presents the estimates obtained using the IV Tobit model.
Table 5. IV Tobit specification for lee1995semi
@@
Other household income $0.973^{***}$
@@
The first column presents the first-stage estimates indicating the relevance of the potential $IV$. The second column presents the structural equation reflecting the effect the treatment variable has on the outcome equation. The parameter $\rho$ shows evidence of no correlation between the unobservables driving the “other” household income and hours worked after controlling for covariates. The estimated coefficient of interest indicates that other household income negatively affects the wife’s labor supply after conditional on the household characteristics. In particular, for women working positive hours, an increase of one thousand dollars in household income from other sources is estimated to reduce hours worked by 9.7 hours per year. The direction of the impacts at the intensive margin of hours worked follows intuitive patterns and is qualitatively similar to those in lee1995semi. However, the model is rejected at conventional significance levels when we test for the IV-Tobit model’s validity, ($\hat{\theta}_{0.99}=0.1335>0$, $\hat{\theta}_{0.95}=0.1368>0$, and $\hat{\theta}_{0.90}=0.1385>0$). This indicates that the assumptions underlying the IV-Tobit model are not compatible with the empirical distribution of the data, and caution is needed when relying on the results.
@@
$\alpha_{1}$ -0.973 -0.419
(0.373) (0.125)
Note: Standard errors in parentheses.
The first column of Table 6 repeats the estimate for $\alpha_{1}$ from Table 5. The second column reports the estimated lower bound for $\alpha_{1}$, obtained under Assumption 6 and latent linear index only based on the nonparametric estimate of the average derivative of the conditional expectation for yearly hours worked with respect to the household income from other sources. As described in Remark 3, we opt to use the estimate for the average derivatives due to instability of the estimated derivatives across the different values of covariates and treatment. This is a conservative approach regarding the bounds for $\alpha$. Inference for the average derivative can be obtained by bootstrapping. Even after relaxing the normality of errors and treatment exogeneity, the lower bound for the size of the effect of having higher household income from sources other than the wife’s labor on their labor supply indicates an effect larger than -4.19 hours worked per year for married women. Hence, we can rule out annual reductions of more than 4.2 hours in female labor supply for every one thousand dollars in other household income, but cannot reject that the effect is zero or positive.